• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 13:10:29 on February22,2002

    Entry number 3542

    keyword=DAQ problems

    Ting, Bob, and 2 guys from DAQ group (Vardan and Ed J.) investigating DAQ
    problems. First we found a bad 1877S in slot 10. Replacing that made ROC4
    work again. Next, Ting wanted to investigate why the fastbus busy signal from
    the Left spectrometer was sometimes a very long signal (roughly 1 in 10 signals
    fluctuate to ~100 msec, the normal is ~30 microsec). We traced this to the back-
    plane of ROC2, the VDC crate. The decision for now is to leave this busy signal
    out of the ganged busy. It was decided not to investigate this further, and to close
    up the halla and resume data taking. It will be decided later if to run in buffered or
    unbuffered mode.